A firm has decided to sell equipment that they purchased five years ago for $10,000. The firm depreciated 50% of the equipments value. The firm is planning on selling the equipment for $6,296. If the tax rate is 25% what is the net sale price (after tax salvage value)?
Answer #1

Book value as on today = 10000 * 50% = 5,000

Profit on Sale 6,296 - 5,000 = $ 1,296

Tax on Profits = profits * taxes  

= 1296 * 25%

= 324

Net Sale Price( Net of Taxes) = Sale Price - taxes

= 6,296 - 324

=$ 5,972

The Net Sale Price is $ 5,972

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.